OpenStudy (anonymous):

What is the solution set when you use the comparison method on "2p + q=1" and "9p + 3q + 3 = 0"

OpenStudy (austinl):

I'm not sure what the comparison method is. I will have to do some research.

OpenStudy (amistre64):

i think its another name for elimination

OpenStudy (amistre64):

either that or a trial and error system ...

OpenStudy (austinl):

The only two methods that I am familiar with are elimination and substitution. Trial and error I don't typically reccommend.

OpenStudy (amistre64):

trial and error is the backbone of mathematics :)

OpenStudy (amistre64):

q= 1-2p q= -1-3p when does 1-2p = -1-3p?

terenzreignz (terenzreignz):

^That is the comparison method... Solve for one variable in both equations, then 'compare' their respective expressions... If that made any sense :D